# Token-Driven GammaTune: Adaptive Calibration for Enchanced Speculative Decoding

### v1

@@ Abstract
Token-Driven GammaTune: Adaptive Calibration for Enchanced Speculative Decoding
Aayush Gautam, Susav Shrestha, Narasimha Annapareddy Department of Electrical and Computer Engineering Texas A&M University, College Station, TX {aayushgautam, sls7161, reddy}@tamu.edu
@@ 1 Introduction
Large language models (LLMs) have become integral to various NLP domains, including information retrieval, conversational AI, and document summarization, driving advancements in language understanding and generation. As these applications demand real-time responsiveness and scalability, optimizing LLM inference is critical for enhancing efficiency and enabling the deployment of high-performance NLP systems Devlin et al. (2019); Shrestha et al. (2024); Ouyang et al. (2022).
@@ 2 Related Work
Since the introduction of speculative decoding by Leviathan et al. and Chen et al., numerous efforts have been made to enhance its efficiency Liu et al. (2024); Xiong et al. (2024); Sun et al. (2024); Yang et al. (2024); Narasimhan et al. (2024); Xia et al. (2024); Wertheimer et al. (2024); Zhou et al. (2024); He et al. (2024); Bhendawade et al. (2024).
@@ 6 Results
GammaTune+ (Ours) $\mathbf{1.28\pm 0.02\times}$ $\mathbf{1.13\pm 0.02\times}$ $1.06\pm\mathbf{0.02}\times$ $\mathbf{1.18\pm 0.05\times}$ $\mathbf{1.16\pm 0.03\times}$

### v3

# Token-Driven GammaTune: Adaptive Calibration for Enhanced Speculative Decoding

@@ Abstract
Token-Driven GammaTune: Adaptive Calibration for Enhanced Speculative Decoding
Aayush Gautam, Susav Shresthathanks: Equal Contribution., Narasimha Reddy Department of Electrical and Computer Engineering Texas A&M University, College Station, TX {aayushgautam, sls7161, reddy}@tamu.edu
@@ 1 Introduction
Large language models (LLMs) have become integral to various NLP domains, including information retrieval, conversational AI, and document summarization, driving advancements in language understanding and generation Devlin et al. (2019); Shrestha et al. (2024, 2025a). As these applications demand real-time responsiveness and scalability, optimizing LLM inference is critical for enhancing efficiency and enabling the deployment of high-performance NLP systems Ouyang et al. (2022); Shrestha et al. (2025b).
@@ 2 Related Work
Since the introduction of speculative decoding by Leviathan et al. and Chen et al., numerous efforts have been made to enhance its efficiency Liu et al. (2024); Xiong et al. (2024); Sun et al. (2024); Yang et al. (2024); Narasimhan et al. (2024); Wertheimer et al. (2024); Zhou et al. (2024); He et al. (2024); Bhendawade et al. (2024).
@@ 6 Results
GammaTune+ (Ours) $\mathbf{1.28\pm 0.02}\times$ $\mathbf{1.13\pm 0.02}\times$ 1.06±0.02×1.06$\pm$ 0.02$\times$1.06 ± 0.02 × $\mathbf{1.18\pm 0.05}\times$ $\mathbf{1.16\pm 0.03}\times$
@@ 8 Limitations
Acknowledgments
This research was supported by the National Science Foundation under Grant No. 2203033.